The following table was made by a student
R: 1 2 3 4 5
1/R: 0.1 0.05 0.033 0.025 0.02

What is the major error in the above table?
1
Expert's answer
2018-11-05T15:44:09-0500

The errors are:

〖∆R_1=R(∆R/R)〗_1=0.1(1)=0.1

〖∆R_2=R(∆R/R)〗_2=0.05(2)=0.1

〖∆R_3=R(∆R/R)〗_3=0.033(3)=0.099

〖∆R_4=R(∆R/R)〗_4=0.025(4)=0.1

〖∆R_5=R(∆R/R)〗_5=0.02(5)=0.1

The error is ∆R/R= 0.033.